OpenStudy (anonymous):

which ratio forms a proportion with 9 over 15 (fraction) a.6 over 10 (fractions) b.16 over 21 (fractions) c.36 over 50 (fractions) d.45 over 70 (Fractions)

OpenStudy (anonymous):

a) cause \[\frac{ 9 }{ 15 } = \frac{ 3*3 }{ 5*3 } = \frac{ 3 }{ 5 } = \frac{ 3*2 }{ 5*2 } = \frac{ 6 }{ 10 }\]

OpenStudy (anonymous):

thank you! :D

OpenStudy (anonymous):

No problem, glad I could help :)